OpenStudy (anonymous):

show me how to simplify 37+8x=4(7-x)

OpenStudy (anonymous):

First we must distribute the 4 into the brackets on the right side. This makes the equations: 37+8x=28-4x. Next we try and get the variable by itself on one side. We will subtract 28 from both sides. This gives us: 9+8x=-4x. Next we subtract 8x from both sides giving us: 9=-12x. Next divide both sides by -12. This gives us 9/12=x. This can further be reduced to -3/4=x